### v2

# Moving Aggregate Modified Autoregressive Copula-Based Time Series Models (MAGMAR-Copulas)

Abstract
Copula-based time series models implicitly assume a finite Markov order. In reality a time series may not follow the Markov property. We modify the copula-based time series models by introducing a moving aggregate (MAG) part into the model updating equation. The functional form of the MAG-part is given as the inverse of a conditional copula. The resulting MAG-modified Autoregressive Copula-Based Time Series model (MAGMAR-Copula) is discussed in detail and distributional properties are derived in a D-vine framework. The model nests the classical ARMA model and can be interpreted as a non-linear generalization of the ARMA-model. The modeling performance is evaluated by modeling US inflation. Our model is competitive with benchmark models in terms of information criteria.
keywords:
Copula , Copula-based time series , Dependence Modeling , Non-linear time series , Persistency , Vine Copula
journal: Bernoulli
1 Introduction
Copula-based time series models are a class of probabilistic autoregressive time series models which allow for flexible temporal dependence modeling as well as flexible marginal modeling. The idea amounts to modeling (possibly non-linear) autoregressive temporal dependencies by a copula while modeling marginal properties independently. Given a univariate, strictly stationary time series $\{X_{t}\}_{t=1}^{T}$, the model decomposes the joint distribution, $F_{X_{t},\ldots,X_{t-p}}$ of $(p+1)$ consecutive random variables of the time series into their stationary marginal distribution, $F_{X}:=F_{X_{t}}=F_{X_{t-1}}=\ldots=F_{X_{t-p}}$, and their copula $C=F_{F_{X}(X_{t}),\ldots,F_{X}(X_{t-p})}$. The marginal distribution, $F_{X}$, accounts for marginal properties such as mean, variance, skewness etc., while the copula accounts for the temporal dependence structure. The general idea was formulated by Darsow et al. (1992) for Markov chains and further investigated in (amongst others) Ibragimov (2005). A short review is given in Patton (2009). Chen and Fan (2006) and Beare (2010) developed the idea for time series. Simard and Rémillard (2015) explore multivariate copula-based time series forecasting with standard copulas while Beare and Seo (2015), Brechmann and Czado (2015) and Nagler et al. (2022) explore multivariate time series modeling using vine-copulas.
These models can be used to model non-linearities in temporal dependencies such as asymmetry; see Müller and Reuber (2022) for irradiation data and McNeil and Bladt (2022) for US inflation. For modeling of heavy-tailedness in the temporal dependence, see Loaiza-Maya et al. (2018) and Pappert and Arsova (2023). Linear temporal dependencies may also be modeled simply, by employing the normal copula. When choosing the copula to be the normal copula and the marginal distribution to be the normal distribution, the linear AR model is recovered (Chen and Fan, 2006, Smith et al., 2010). Thus the model can be understood as a generalization to the AR model which permits broader temporal dependencies and marginal distributions.
One caveat of the copula-based time series models is that, implicitly, the process is ascribed to be a Markov process of order $p$, i.e. the true data-generating process is of finite autoregressive order $p$. This means that the variable $X_{t}$ may depend explicitly on a finite number of past values $X_{t-1},\ldots,X_{t-p}$ only. However, in reality the Markov assumption may fail depending on the time series. The time series that is to be modeled may display persistent behavior i.e. non cut-off autodependence. In principle it is possible to extend the copula-based time series model to arbitrary Markov order, however it becomes unfeasible as the number of parameters increases indefinitely. Hence, we think that there is a need for a copula-based time series model with infinite order partial dependence. To this end, McNeil and Bladt (2022) propose a possible solution: They consider the copula-based time series model with an infinite autoregressive order. Hence, for each autoregressive order, there is a partial copula describing the corresponding temporal dependence. Handling such a model is, without further manipulation, impossible because it has infinitely many parameters. However, when reparametrizing it in a clever way, the model becomes manageable. The reparametrization proposed by the authors is based on linear ARMA models. The caveat of their model is that it is cumbersome to incorporate copulas with more than one parameter, e.g. the $t$-copula. In contrast, our model can accommodate general copulas by following a conceptually different approach. The approach by McNeil and Bladt (2022) is reviewed in Section 2.4.
Another approach to generalize copula-based time series models to allow for infinite autodependence is given in Chapter 3.14 of Joe (2014). Therein, a model, which is quite similar to the MAGMAR-copula model, is introduced. In particular, what we call ’moving aggregate’, he calls a ’$q$-dependent’ extension, and a copula extension to ARMA processes is named ’Mixed Markov and $q$-dependent’. Though similar, the models are not the same; while our model may be interpreted as feeding a non-linear AR model with noise filtered by a moving aggregate, the model by Joe feeds a non-linear AR-process to a non-linear $q$-dependent filter. We note that the MAGMAR-copula model has been developed independently. We discuss the similarities and differences between the two models in Sect. 2.5.
Now we turn our attention to our proposed model. The problem of limited feasible autoregressive orders in time series models also occures in the AR($p$) and ARCH($p$) models. The solution in the AR($p$) and ARCH($p$) models is to introduce moving average terms, resulting the the ARMA and GARCH models, respectively. (See Engle (1982), Bollerslev (1986), Bollerslev et al. (1994), Francq and Zakoian (2019) and Hamilton (2020)). By inverting the corresponding linear filter, AR($\infty$), respectively ARCH($\infty$) representations can be derived. Hence the time series can be understood as having no Markov restriction. This results in the time series models being more persistent in the mean in the case of ARMA and more persistent in the variance in the case of GARCH. The latter can be interpreted as the model being able to capture volatility clusters (Bollerslev et al., 1994).
Inspired by this approach, we modify the copula-based time series model by incorporating a moving aggregate (MAG) part in the updating equation. The MAG part is a non-linear function taking the current and the past $q$ innovations as arguments. Formally it is a mapping from $\mathbb{R}^{q+1}$ to $\mathbb{R}$. The functional form of the MAG-function is given by an inverse conditional copula. We call it ’moving aggregate’, because, similar to the moving average, it aggregates the current and past $q$ innovations but it is no linear aggregation, hence ’aggregate’ instead of ’average’. Analogous to the classical ARMA and GARCH models, the MAG-part permits infinite partial dependence. With this, we mean that the partial copula (see Spanhel and Kurz, 2016) of $X_{t}$ and $X_{t-j}$ is different from the independence copula for all $j,t\in\mathbb{Z}$. We call the resulting model MAG-Modified Autoregressive Copula-Based Time Series Model, or in short: MAGMAR-Copula. In this paper we present the model and show that the classical copula-based time series model and the classical linear ARMA model are nested within the proposed model. We proceed to derive its distributional properties in a D-vine framework (see Aas et al. (2009), Czado (2010), Czado and Nagler (2022) and Sect. 2.1 for D-vine copulas). AR($\infty$) and MAG($\infty$) representations are derived. Building upon the MAG($\infty$) representation, it is proven that the MAGMAR-Copula model, where the AR-copula is the normal copula, is stationary when excluding perfect comonotonicity and perfect countermonotonicity. Furthermore, we derive the conditional distributions of $X_{t}|\mathcal{F}_{t-1}$, where $\mathcal{F}_{t-1}$ denotes the $\sigma$-field of all information up to and including $(t-1)$. Based on conditional independence, the likelihood function is derived. Similar to ARMA and GARCH models, the likelihood has to be calculated in an iterative scheme.
Last, we apply the MAGMAR-Copula model to quarterly US inflation. The parameters are estimated by maximum likelihood estimation. We compare the modeling performance of the MAGMAR-Copula model to that of the copula-ARMA model of McNeil and Bladt (2022) and to linear modeling (where normal copulas as temporal dependence structure). We find that our model performs well in this setting and is competitive with the model from McNeil and Bladt (2022). In line with their results, we find that modeling asymmetric temporal dependence structures (by incorporating the gumbel copula) as well as infinite partial dependence is beneficial. Furthermore, we find that incorporating heavy-tailed temporal dependence structures (i.e. by incorporating the $t$-copula) enhances the performance, indicating not only asymmetric but also heavy-tailed temporal dependencies in US inflation.
In the next section, Sect. 2, the copula-based time series model is introduced and reviewed. We focus on the model class where the copula is modeled as a D-Vine copula. Sect. 2.4 reviews the model for time series with infinite partial copula dependence by McNeil and Bladt (2022). In Sect. 3 the Moving Aggregate Modified Autoregressive Copula-Based Time Series Model (MAGMAR-Copula) is introduced and explored. The MAGMAR-Copula model is applied to US Inflation in Sect. 4. Sect. 5 concludes the work and gives an outlook to future research.
2 Copula-based Time Series Models
In this section the notion of copula-based time series models is introduced. First, we give a short introduction to copulas. Then general ideas of copula-based time series is developed for the Markov(1) case and then extended to arbitrary order. Last, we discuss the copula-ARMA model, which permits infinite partial dependence, proposed by McNeil and Bladt (2022).
2.1 Copulas and D-vine Copulas
In this section we give a short introduction to copulas and the main ideas to multivariate modeling with copulas. For detailed and comprehensive presentations we refer to standard works: Joe (2014) and Nelsen (2006).
Consider $p$ continuous random variables $X_{1},\ldots,X_{p}$. Let their joint distribution be given by $F_{X_{1},\ldots,X_{p}}$ and their marginal distributions by $F_{X_{1}},\ldots,F_{X_{p}}$, respectively. The general idea of multivariate copula modeling amounts to decomposing the joint distribution into marginal distributions and a copula,
$F_{X_{1},\ldots,X_{p}}(x_{1},\ldots,x_{p})=C[F_{X_{1}}(x_{1}),\ldots,F_{X_{p}}(x_{p})].$ (1)
Sklar’s theorem states that there exists a copula that fulfills the statement from Eq. 1, which is unique if the random variables are continuous (Sklar, 1959, Joe, 2014). The copula $C$ is the joint distribution of probability integral transformed (PIT) and hence uniform random variables $F_{X_{1}}(X_{1}),\ldots,F_{X_{p}}(X_{p})$ (see Angus, 1994, for the PIT), $F_{F_{X_{1}}(X_{1}),\ldots,F_{X_{p}}(X_{p})}=C$. The uniformity of $U_{i}:=F_{X_{i}}(X_{i}),\ i\in\{1,\ldots,p\}$ allows the copula to be interpreted as the dependence structure of the random variables $X_{1},\ldots,X_{p}$. An important notion for visualization, estimation and for D-vine copulas is the copula density. The copula density of a copula $C$ is given as (Joe, 2014)
$c[u_{1},\ldots,u_{p}]=\frac{\partial^{p}C[u_{1},\ldots,u_{p}]}{\partial u_{1}\ldots\partial u_{p}}.$ (2)
Since the copula density is given by the partial derivative of the respective copula, Eq. 2 is only valid if the copula is differentiable. In applications, the copula and the marginal distributions have to chosen appropriately. The normal, t, gumbel and independence copula are described in G. There are other important copulas that deserve being mentioned here, however, we restrict the exposition to the copulas that are used in this work for brevity. Again we refer to Joe (2014) and Nelsen (2006) for more examples of copulas.
Now we introduce the notion of D-vine copulas. D-vine copulas are a special case of pair-copula constructions. A $p$-variate copula is decomposed into pair-copulas, i.e. bivariate copulas (Bedford and Cooke, 2001, 2002, Aas et al., 2009, Czado, 2010, Czado and Nagler, 2022). The decomposition is based on the law of total probability and is not unique. The D-vine decomposition amounts to a special ordering of the variables in the decomposition. The decomposition in terms of the $p$-variate copula density is given as (Aas et al., 2009),
$c[u_{1},\ldots,u_{p}]=$ $\prod_{j=1}^{p-1}\prod_{i=1}^{n-j}c_{i,i+j|i+1,\ldots,i+j-1}\Big{(}F(u_{i}|u_{i+1},\ldots,u_{i+j-1}),$
$F(u_{i+j}|u_{i+1},\ldots,u_{i+j-1})\Big{)},$ (3)
where $c_{i,i+j|i+1,\ldots,i+j-1}$ is the conditional copula density of $U_{i}$ and $U_{i+j}$ given the intermediate variables $U_{i+1},\ldots,U_{i+j-1}$. $F(u_{i}|u_{i+1},\ldots,u_{i+j-1})$ is the conditional distribution function of $U_{i}$ given $U_{i+1},\ldots,U_{i+j-1}$ evaluated at $u_{i}$ and conditioning variables $u_{i+1},\ldots,u_{i+j-1}$. Since the joint distribution of $U_{i},U_{i+1},\ldots,U_{i+j-1}$ is a copula, $F(u_{i}|u_{i+1},\ldots,u_{i+j-1})$ may also be understood as a conditional copula. The decomposition in Eq. 3 is called ’D-vine’ for ’drawable vine’ because the variables are ordered in a path like manner (Czado and Nagler, 2022). The vine copula approach allows for more flexibility when modeling multivariate dependencies because each bivariate dependence can be specified separately.
2.2 Copula-based Time Series with Autoregressive Order of One
Consider a strictly stationary time series $\{X_{t}\}_{t=1}^{T}$. A time series being Markov(1) means that $\mathbb{P}(X_{t}\leq a|X_{t-1}=x_{t-1},\ldots,X_{1}=x_{1})=\mathbb{P}(X_{t}\leq a|X_{t-1}=x_{t-1}),\ t>1$. Accordingly a time series of Markov order $p$ fulfills $\mathbb{P}(X_{t}\leq a|X_{t-1}=x_{t-1},\ldots,X_{1}=x_{1})=\mathbb{P}(X_{t}\leq a|X_{t-1}=x_{t-1},\ldots,X_{t-p}=x_{t-p}),\ t>p$ (Ibe, 2013). First, we consider the case where $p=1$. Afterwards, we consider the case for arbitrary $p$. For $p=1$, the time series is Markov($1$) and characterized by the bivariate joint distribution $F_{X_{t},X_{t-1}}$ of consecutive random variables, $X_{t}$ and $X_{t-1}$. The idea is to decompose the joint distribution into a copula and marginal distributions. By Sklar’s Theorem there always exists a copula $C:=F_{F_{X}(X_{t}),F_{X}(X_{t-1})}$ such that the joint distribution can be decomposed into marginal distributions $F_{X_{t}}$ and $F_{X_{t-1}}$ and the copula (Sklar, 1959, Joe, 2014). If the random variables $X_{t}$ and $X_{t-1}$ are continuous, the decomposition is unique. Due to strict stationarity, $F_{X_{t}}=F_{X_{t-1}}=:F_{X}$. The decomposition may be written as
$F_{X_{t},X_{t-1}}(a,b)=C[F_{X}(a),F_{X}(b)].$
Hence, the joint distribution is specified by a marginal distribution $F_{X}$ and the copula $C$. Note that the copula is the joint distribution of the probability-integral transformed random variables $U_{t}:=F_{X}(X_{t})$ and $U_{t-1}:=F_{X}(X_{t-1})$. The copula may be parametrized by dependence parameter(s) $\phi$, so $C=C_{\phi}$. We introduce this notation to emphasize the analogy to the linear AR model and later on, in Sect. 3 to the linear ARMA model. Then, analogous to the linear AR model, the Markov($1$) model can be formulated in a time series model equation as
$X_{t}=F^{-1}_{X}\left(C_{\phi}^{-1}[w_{t}|F_{X}(X_{t-1})]\right),\ \ \ w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1).$
We note that this or similar formulations are also used in other works under different names such as ’updating equation’ in McNeil and Bladt (2022) or in descriptions on how to sample from copulas in Aas et al. (2009). The calculation of the new realization, $x_{t}$, amounts to first sampling a realization from a uniform distribution, $w_{t}$ and then quantile transforming $w_{t}$ with the conditional copula obtaining $U_{t}:=C_{\phi}^{-1}[w_{t}|F_{X}(X_{t-1})]$. Last, $U_{t}$ is again quantile transformed with the marginal distribution quantile function $F_{X}^{-1}$. Using this procedure, it is guaranteed that the dependence structure of $X_{t}$ and $X_{t-1}$ is given by $C_{\phi}$ and that the marginal distribution is given by $F_{X}$. This model can be regarded as a generalization of the gaussian AR(1)-model. This can be seen by choosing $C_{\phi}$ as the bivariate normal copula $C[u_{1},u_{2}]=\Phi_{\phi}(\varphi^{-1}(u_{1}),\varphi^{-1}(u_{2}))$ (see e.g. Joe (2014) for details about the normal copula or Sect. 2.1) and the marginal distribution $F_{X}$ as the standard normal distribution. Although this is a well known fact in the literature, see e.g. Chen and Fan (2006) or Smith et al. (2010), we give the derivation in Appendix A for completeness. In the following only the pseudo-observation time series, $\{U_{t}\}_{t=1}^{T},\ U_{t}:=F_{X}(X_{t})$, is considered. The actual time series can be recovered by quantile transforming every pseudo-obersvation by $F_{X}^{-1}$. The dynamics of the time series are solely determined by the relation between the pseudo-observations. The model equation of the pseudo-observations is given as,
$U_{t}=C^{-1}_{\phi}[w_{t}|U_{t-1}],\ \ w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1).$ (4)
The non-linearity of the model is two-fold. First, the relation between random variables $U_{t}$ and $U_{t-1}$ is a non-linear function. Second, the innovations $w_{t}$ enter in a non-additive way. What at first sight may seem strange is actually necessary. Choosing the model equation this way is the only possibility to ensure that $U_{t},U_{t-1},\ldots$ are uniformly distributed and that $C_{\phi}$ is the copula of consecutive variables. The uniformity of the unconditional distribution of all $U_{t},\ t\in\mathbb{Z}$ is due the properties of inverse conditional copulas: Given $U_{1},U_{2}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1)$ it follows that $C^{-1}_{\phi}[U_{1}|U_{2}]\sim U(0,1)$. This implies that all $U_{t}$ are distributed identically. From this it follows that the time series is stationary, because the transitions in the updating equation, Eq. 4, are homogeneous.
2.3 Copula-based Time Series models with an arbitrary autoregressive order
The model can be generalized in the case where $\{X_{t}\}_{t\in\mathbb{Z}}$ is a strictly stationary Markov($p$) time series. The behavior of the time series is determined by the joint distribution of $p+1$ consecutive random variables $X_{t},X_{t-1},\ldots,X_{t-p}$. As before, decomposing the joint distribution using Sklar’s Theorem yields
$F_{X_{t},\ldots,X_{t-p}}(x_{t},\ldots,x_{t-p})=C[F_{X}(x_{t}),\ldots,F_{X}(x_{t-p})],$
where again $F_{X}:=F_{X_{t}}=\ldots=F_{X_{t-p}}$ due to strong stationarity and $C$ is the joint distribution of the individually uniformly distributed random variables $(U_{t},\ldots,U_{t-p})$ where $U_{t}:=F_{X}(X_{t})\ \forall t\in\mathbb{Z}$. Considering the time series of pseudo-observations $\{U_{t}\}_{t=1}^{T}$ and parametrizing the copula with dependence parameter(s) $\phi$, the time series model equation can be formulated as
$U_{t}=C_{\phi}^{-1}[w_{t}|U_{t-1},\ldots,U_{t-p}],\ \ w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1).$ (5)
Note that the copula $C_{\phi}$ is now a $(p+1)$-dimensional copula. The parameter $\phi$ may be a scalar or a vector, depending on the chosen copula. Choosing a specific copula as $C_{\phi}$ may restrict the model to one dependence structure. In order to account for possibly different forms of dependency and to be flexible in modeling, D-vine copulas (Bedford and Cooke, 2001, 2002, Aas et al., 2009, Czado, 2010, Czado and Nagler, 2022) can be employed for modeling the temporal dependence as is also done in Nagler et al. (2022) and Smith et al. (2010). In this approach the joint density of $(p+1)$ consecutive pseudo-observations is decomposed into a product of bivariate copulas. The decomposition is given by (Smith et al., 2010, McNeil and Bladt, 2022, Nagler et al., 2022)
$c_{t,\ldots,t-p}(u_{t},\ldots,u_{t-p})=$ $\prod_{k=t-p}^{t-1}\prod_{j=k+1}^{t}c_{j-k,j|S_{j-k,j}}\Big{[}F_{j-k|S_{j-k,j}}(u_{j-k}),$
$F_{j|S_{j-k|j}}(u_{j})\Big{]},$ (6)
where the shorthand notation $c_{t,\ldots,t-p}=c_{U_{t},\ldots,U_{t-p}}$ is used. Furthermore, $S_{j-k,j}=\{j-k+1,\ldots,j-1\}$. As in the cited literature, only the simplified D-vine is considered, dropping the conditioning in the bivariate copula densities, $c_{j-k,j|S_{j-k,j}}=c_{j-k,j}$. In the case where $c_{j-k,j}=c_{i-k,i}:=c_{k}\ \forall j,i\in\{k+1,\ldots,T\}$, the joint density of all random variables $\{U_{t}\}_{t=1}^{T}$ admits translational invariance. Nagler et al. (2022) show that such processes are stationary. They also may be called s-vine (’s’ for stationary) (McNeil and Bladt, 2022). Furthermore, the bivariate copulas, $C_{j-k,j}$ may be parametrized by parameters $\phi_{1},\ldots,\phi_{p}$, where $\phi_{1}$ governs the dependence of $U_{t-1}$ and $U_{t}$ $\forall t\in\mathbb{Z}$, $\phi_{2}$ the dependence of $U_{t}|U_{t-1}$ and $U_{t-2}|U_{t-1}$ and so on. Note that depending on the chosen copula, each $\phi_{i},i\in\{1,\ldots,p\}$ may also be a vector containing more than one parameter.
It is common to denote conditional distributions of pseudo observations by the $h$-function (Aas et al., 2009, Czado, 2010, Czado and Nagler, 2022). The $h$-functions associated to a copula $C$ are defined as
$h^{1}(u_{1},u_{2})=\frac{\partial C[u_{1},u_{2}]}{\partial u_{1}},$
$h^{2}(u_{1},u_{2})=\frac{\partial C[u_{1},u_{2}]}{\partial u_{2}}.$ (7)
If no superscript is given, then conditioning on the second variable is implied. The conditional distribution $F_{U_{t}|U_{t-1}}$ may be rewritten as $F_{U_{t}|U_{t-1}}(u_{t},u_{t-1})=C_{\phi}(u_{t}|u_{t-1})=h^{2}_{\phi_{1}}(u_{t},u_{t-1})$. The same applies to $U_{t-2}|U_{t-1}$: $F_{U_{t-2}|U_{t-1}}(u_{t-2},u_{t-1})=C_{\phi}(u_{t-2}|u_{t-1})=h^{1}_{\phi}(u_{t-1},u_{t-2})$. The $h$-functions can be derived from the associated copula. Their calculation amounts to taking the derivative of copulas. Often the $h$-functions can be derived analytically. (Explicit forms of $h$-functions and their inverses can be found in Smith et al. (2010) for the normal, t, clayton, gumbel and galambos copula.) An advantage of using the $h$-functions is that all conditional distributions can be decomposed into a function of $h$-functions. E.g. the distribution of $(U_{t}|U_{t-1},U_{t-2})$ which is needed for the partial distribution of $U_{t}$ and $U_{t-3}$, can be calculated as (Aas et al., 2009, Joe, 2014)
$F_{U_{t}|U_{t-1},U_{t-2}}(u_{t}|u_{t-1},u_{t-2})=h_{\phi_{2}}^{2}\big{(}h^{2}_{\phi_{1}}(u_{t},u_{t-1}),h^{1}_{\phi_{1}}(u_{t-1},u_{t-2})\big{)}.$
The notion of decomposing a conditional distribution into a function of $h$-functions is named Rosenblatt (forward/backward) function (McNeil and Bladt, 2022). The decomposition of $U_{t}|U_{t-1},\ldots,U_{t-p}$ can be called Rosenblatt backward and the decomposition of $U_{t-p}|U_{t-p+1},\ldots,U_{t}$ Rosenblatt forward function (inspired by the Rosenblatt transform from Rosenblatt (1952)). For $p=1$ the Rosenblatt functions are just the corresponding $h$-functions, $R_{1}^{1}=h_{1}^{1}$, $R_{1}^{2}=h_{1}^{2}$. McNeil and Bladt (2022) define the Rosenblatt forward and backward functions for $p>1$, $R_{p}:(0,1)^{p}\times(0,1)\rightarrow(0,1)$, recursively as follows:
$R_{p}^{1}(\mathbf{u},x)=$ $h_{p}^{1}\left(R_{p-1}^{2}(\mathbf{u}_{-1},u_{1}),R_{p-1}^{1}(\mathbf{u}_{-1},x)\right),$
$R_{p}^{2}(\mathbf{u},x)=$ $h_{p}^{2}\left(R_{p-1}^{2}(\mathbf{u}_{-p},x),R_{p-1}^{1}(\mathbf{u}_{-p},u_{p})\right).$ (8)
Here $\mathbf{u}_{-j}$ denotes $\mathbf{u}_{-j}=(u_{1},\ldots,u_{j-1},u_{j+1},\ldots,u_{k})$. The second argument of the Rosenblatt functions, $x$, is the actual variable that the Rosenblatt functions are evaluated at. The conditioning variables are contained in the vector $\mathbf{u}$. Using the newly introduced notation, the D-vine copula density (Eq. 6) can be rewritten as (McNeil and Bladt, 2022),
$c_{t,\ldots,t-p}(u_{t},\ldots,u_{t-p})=\prod_{k=t-p}^{t-1}\prod_{j=k+1}^{t}c_{j-k,j}\Big{[}$ $R^{1}_{k-1}((u_{j-k+1},\ldots,u_{j-1}),u_{j}),$
$R^{2}_{k-1}((u_{j-k+1},\ldots,u_{j-1}),u_{j-k})\Big{]}.$
In order to rewrite the time series updating equation (Eq. 5) in the D-vine framework, the inverse Rosenblatt transforms can be utilized:
$U_{t}=R^{-1}_{p}((U_{t-p},\ldots,U_{t-1}),w_{t}),\ w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1).$ (9)
Here $R_{p}$ is the Rosenblatt function corresponding to the $(p+1)$-dimensional copula $C_{\phi}$ from Eq. 5. In this notation, the Markov($p$) restriction becomes obvious. The model can also be understood as a generalized AR($p$)-model. In this case, however, the function connecting the past variables, $U_{t-1},\ldots,U_{t-p}$, innovations $w_{t}$ and the current observations $U_{t}$ is not necessarily linear. For the case where $p=1$, the updating equation becomes
$U_{t}=h_{\phi}^{-1}(w_{t},U_{t-1}),\ w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1).$ (10)
Here $h_{\phi}$ is the $h$-function associated to the 2-dimensional copula $C_{\phi}$ from Eq. 4. Next, we are going to discuss the extension of McNeil and Bladt (2022) to allow for infinite partial dependence.
2.4 Review of the extension by McNeil & Bladt
McNeil and Bladt (2022) allow for infinite partial dependence in copula-based time series models by letting the autoregressive order, $p$, be infinite and reparametrizing the joint density with a finite amount of parameters. The parameters are chosen in such a way that they represent a classical ARMA($p,q$) process with finite $p$ and $q$. To model a time series, the following steps are employed. First, the autoregressive and the moving average order, $p$ and $q$, respectively, are determined by minimizing model selection criteria (e.g. AIC). Next, the partial autocorrelations of the chosen ARMA($p,q$) process are parametrized in terms of the autoregressive and the moving average parameters, $\phi_{1},\ldots,\phi_{p},\theta_{1},\ldots,\theta_{q}$. Using the one-to-one relation between autocorrelations and Kendall’s $\tau$, the Kendall’s partial autocorrelations are obtained. In a third step, the partial copulas in Eq. 6 are reparametrized by Kendall’s $\tau$ and consecutively by the AR and MA parameters, which can be estimated by maximum likelihood.
One disadvantage of their extension is that the Kendall’s $\tau$ reparametrization only works for one-parametric copulas, since only for one-parametric copulas, there is a one-to-one relation. This excludes, for example, the t-copula, which is of particular relevance in practice (see e.g. Loaiza-Maya et al., 2018, Pappert and Arsova, 2023). Reparametrization for copulas with more than one parameter is more cumbersome.
2.5 Review of the extension by Joe
Now we want to briefly introduce the model of Joe (2014), which is to the best of our knowledge the earliest copula extension of the ARMA model. The model is quite similar to the MAGMAR-copula model, although it is not the same. The model with uniform marginals is given by the following updating equations:
$W_{s}=$ $h(\varepsilon_{s};W_{s-1},\ldots,W_{s-p}),\ \ s\in\mathbb{Z},$ (11)
$U_{t}=$ $g(\varepsilon_{t};W_{t-q},\varepsilon_{t-1},\ldots,\varepsilon_{t-q+1}),\ \ t\in\mathbb{Z},$ (12)
where $\varepsilon_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1)$ are the common innovations, $h$ is the conditional quantile function of a $(p+1)$-variate copula and $g$ is the conditional quantile function of a $(q+1)$-variate copula. It is required that the copula corresponding to $h$ is translationally invariant and that the $q$-variate marginal distribution of the copula corresponding to $g$ is the independence copula. The idea is inspired by the Poisson ARMA process of McKenzie (1988): First create a process corresponding to an AR process, Eq. 11, and then feed this process to an MA-like filter together with the common innovations, Eq. 12. In our proposed MAGMAR-copula model, we too attempt to mix AR-type and MA-type copula processes. The difference, however, is that we feed the MA-type process (what we call ’MAG’) into the AR-type filter. In other words, in the copula extension to ARMA models by Joe (2014) the latent process is an AR-type process whereas in the MAGMAR-copula model, the latent process is an MA-like model.
3 MAGMAR-Copulas
In this section, we introduce and discuss the moving aggregate modification to copula-based time series models. For illustration, the autoregressive and moving aggregate orders are set to one. Results for arbitrary orders $p$ and $q$ can be derived in a similar fashion; Some results with arbitrary orders are given in the Appendix.
3.1 Introduction to the model
Our proposed approach enabling the model to overcome the Markov restriction amounts to modifying the autoregressive updating equation, Eq. 10, by a moving aggregate part. The moving aggregate part is, similar to classical, linear ARMA processes, a function of the current and past innovations. It is given as an inverse conditional copula, which we parametrize with $\theta$. So instead of $w_{t}$ in Eq. 10, we have $h_{\theta}^{-1}(w_{t},w_{t-1})$. We call the model Moving Aggregate Modified Autoregressive Copula-Based Time Series Model (MAGMAR-Copula model). In the following exposition, we assume that 1) all copulas are differentiable, 2) that they are within the Fréchet-Hoeffding bounds (see Joe, 2014, Chapter 2.9) and 3) that the inverses of $h$-functions associated to all used copulas exist.
Definition 1 (MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula model).
Let $h^{-1}_{\phi}:(0,1)^{2}\rightarrow(0,1)$ and $h_{\theta}^{-1}:(0,1)^{2}\rightarrow(0,1)$ be inverse $h$-functions associated to copulas $C_{\phi}$ and $C_{\theta}$, which are parametrized by parameters $\phi$ and $\theta$, respectively. The time series $\{U_{t}\}_{t\in\mathbb{Z}}$ following the updating equation
$U_{t}=h_{\phi}^{-1}(h_{\theta}^{-1}(w_{t},w_{t-1}),U_{t-1}),\ \ \ w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1)$ (13)
is a MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula time series. The corresponding model is the MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula model. The copula $C_{\phi}$ is the AR-Copula, while $C_{\theta}$ is the MAG-Copula. $h_{\theta}^{-1}(w_{t},w_{t-1})$ is the MAG-part.
The $h$-functions associated to the AR and MAG-copula can be derived from the respective copulas using Eq. 7. The family of the copulas (e.g. normal, t, gumbel, etc.) has to be chosen. Parameters $\phi$ and $\theta$ have to be set/estimated. We note that the MAG-part, that is used here, was also derived in McNeil and Bladt (2022) as the ’causal representation’ for the classical copula-based time series models. The new contribution is that we use this representation to construct the new MAGMAR-Copula model.
Before giving the relations to other models and deriving distributional properties, we proceed to discuss implications of the model equation and further explain the model. First, we note that $C_{\theta}$ is not the copula of $w_{t}$ and $w_{t-1}$. The innovations $w_{t}$ and $w_{t-1}$ are independent and their copula is the independence copula, $F_{w_{t},w_{t-1}}=\Pi$. The expression $h_{\theta}(w_{t},w_{t-1})$ is to be understood as a function of two random variables $w_{t}$ and $w_{t-1}$, just as $\varepsilon_{t}-\theta\varepsilon_{t-1}$ in the classical ARMA model is a function of two random variables $\varepsilon_{t}$ and $\varepsilon_{t-1}$. Second, The copula $C_{\phi}$ is also not the copula of $U_{t}$ and $U_{t-1}$, however it is connected to the copula of the two random variables. This is analogous to the classical ARMA(1,1) model, where the partial autocorrelation of $X_{t-1}$ and $X_{t}$ is not solely determined by the AR(1) parameter but also by the MA(1) parameter. (The influence of the MAG-Copula and the AR-Copula to the dependence of $U_{t}$ and $U_{t-1}$ becomes clear when we derive the joint conditional distribution of $U_{t}$ and $U_{t-1}$ given $\mathcal{F}_{t-2}$.) The expression $h_{\phi}(h_{\theta}^{-1}(w_{t},w_{t-1}),U_{t-1})$ may also be understood as a function of random variables. To verify that the MAGMAR(1,1)11(1,1)( 1 , 1 ) model equation actually yields the expected behavior to have no Markov restriction i.e. infinite autoregressive order, the model equation can be iterated. To this aim, an explicit form for the innovations is necessary. It can be directly obtained from Eq. 13 as
$w_{t}=h_{\theta}(h_{\phi}(U_{t},U_{t-1}),w_{t-1}).$
Iteratively substituting for $w_{t-1}$ in Eq. 13 yields the AR($\infty$) representation. Utilizing the composition symbol, which acts as $\bigcirc_{i=1}^{N}f_{i}=f_{1}\circ f_{2}\circ\ldots\circ f_{N}$, the representation can be written as
$w_{t}=$ $h_{\theta}(h_{\phi}(U_{t},U_{t-1}),\cdot)\circ h_{\theta}(h_{\phi}(U_{t-1},U_{t-2}),\cdot)\circ\ldots,$
$=$ $\bigcirc_{i=0}^{\infty}h_{\theta}(h_{\phi}(U_{t-i},U_{t-i-1}),\cdot).$ (14)
Hence, in general, each past observation, $U_{t-i},i=1,2,\ldots$ has a non-independent influence on $U_{t}$. The MAG($\infty$) representation can be derived as
$U_{t}=$ $h_{\phi}^{-1}(h_{\theta}^{-1}(w_{t},w_{t-1}),\cdot)\circ h_{\phi}^{-1}(h_{\theta}^{-1}(w_{t-1},w_{t-2}),\cdot)\circ\ldots,$
$=$ $\bigcirc_{i=0}^{\infty}h_{\phi}^{-1}(h_{\theta}^{-1}(w_{t-i},w_{t-i-1}),\cdot).$ (15)
As noted before, there also exists a MAG($\infty$) representation for the case with no moving aggregate extension. The representation was derived in McNeil and Bladt (2022) as ’causal representation’. It is easy to see that for $C_{\theta}=\Pi$, the two representations are equal. This is in line with proposition 2, which proposes that for $C_{\theta}=\Pi$, the MAGMAR(1,1)-Copula model is equal to the classical copula-based time series model. The proposition and the proof are given in Sect. 3.2. In Sect. 3.3.1, it is explicitly shown that the MAG($\infty$) representation of the MAGMAR-Copula time series converges for the case where the AR-copula, $C_{\phi}$, is the normal copula.
Last, we want to introduce the MAGMAR$(p,q)$-Copula model with $p$ autoregressive and $q$ moving aggregate orders. Using the Rosenblatt functions (Eq. 8) It can be formulated as follows:
Definition 2 (MAGMAR$(p,q)$-Copula Model).
Let $R_{\phi_{1},\ldots,\phi_{p}}:(0,1)^{p+1}\rightarrow(0,1)$ be a Rosenblatt function associated to a D-vine copula $C_{\phi_{1},\ldots,\phi_{p}}$, which is parametrized by parameters $\phi_{1},\ldots,\phi_{p}$. Furthermore, let $R^{-1}_{\theta_{1},\ldots,\theta_{q}}:(0,1)^{q+1}\rightarrow(0,1)$ be an inverse Rosenblatt function associated to a D-vine copula $C_{\theta_{1},\ldots,\theta_{q}}$, which is parametrized by parameters $\theta_{1},\ldots,\theta_{q}$. The time series $\{U_{t}\}_{t\in\mathbb{Z}}$ following the updating equation
$R^{1}_{\phi_{1},\ldots,\phi_{p}}((U_{t-p},\ldots,U_{t-1}),U_{t})=(R^{1}_{\theta_{1},\ldots,\theta_{q}})^{-1}((w_{t-q},\ldots,w_{t-1}),w_{t})$ (16)
is a MAGMAR$(p,q)$-time series. The corresponding model is the MAGMAR$(p,q)$-Copula model. The copula $C_{\phi_{1},\ldots,\phi_{p}}$ is the AR-Copula. The copula $C_{\theta_{1},\ldots,\theta_{q}}$ is the MAG-Copula. The LHS of Eq. 16 is the AR-part and the RHS is the MAG-part.
The Rosenblatt functions in Eq. 16 ought to be understood as a composition of $h$-functions (as given in Eq. 8). The AR-Rosenblatt function, $R^{1}_{\phi_{1},\ldots,\phi_{p}}$ is composed of $h$-functions $h_{\phi_{1}},\ldots,h_{\phi_{p}}$ while the MAG-Rosenblatt function is composed of $h$-functions $h_{\theta_{1}},\ldots,h_{\theta_{q}}$. Analogous to to the MAGMAR(1,1)11(1,1)( 1 , 1 ) case, the AR-Rosenblatt function is not to be understood as conditional distribution of $U_{t}$ given $U_{t-1},\ldots,U_{t-p}$, but as a mere function, whose arguments are $U_{t},\ldots,U_{t-p}$. The same holds for the MAG-Rosenblatt function and $w_{t},w_{t-1},\ldots,w_{t-q}$. In the case where the copulas, respectively $h$-functions involved in $R^{1}_{\theta_{1},\ldots,\theta_{t-q}}$ are chosen as the independence copula, the conditional distribution of $U_{t}|U_{t-1},\ldots,U_{t-p}$ is given by $R^{1}_{\phi_{1},\ldots,\phi_{p}}$. The MAG($\infty$)∞ ) and the AR($\infty$) representations of the MAGMAR$(p,q)$-Copula can be derived in a similar fashion to the MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula case. This result is given in Appendix D. Next, we give a conceptual comparison between the MAGMAR-Copula model and the copula-ARMA model by McNeil and Bladt (2022).
3.2 Relation to other models
Now we derive the relation of the proposed MAGMAR-Copula model to the classical copula-based time series model and the classical linear ARMA model. The results are given for the MAGMAR(1,1)11(1,1)( 1 , 1 ) case but can be extended to the MAGMAR(1,1)11(1,1)( 1 , 1 ) case. Furthermore we define the MAG-copula time series and show that it is stationary. The first two proposition show that the classical linear ARMA model as well as the classical copula-based time series model is nested by the MAGMAR-Copula model.
Proposition 1.
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula time series. If the AR-Copula, $C_{\phi}$, and the MAG-Copula, $C_{\theta}$, are chosen as the normal copula, $C_{\phi}=C^{\text{normal}}_{\phi}$, $C_{\theta}=C^{\text{normal}}_{\theta}$ with $|\phi|<1$ and $|\theta|<1$, the quantile transformed time series $\{\varphi^{-1}(U_{t})\}_{t\in\mathbb{Z}}$ is the classical linear ARMA(1,1)11(1,1)( 1 , 1 ) time series model with AR-parameter $\phi$, MA-parameter $\frac{\beta}{(1-\beta^{2})^{\frac{1}{2}}}$ and independent normally distributed innovations with variance $(1-\phi^{2})(1-\theta^{2})$.
Proof.
See Appendix A
Proposition 2.
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula time series. If the MAG-copula is the independence copula, $C_{\theta}=\Pi$, then the classical copula-based time series model (see Sect. 2.1) is recovered.
Proof.
The $h$-function associated to the independence copula is $h_{\Pi}(u_{1},u_{2})=\frac{\partial(u_{1}u_{2})}{\partial u_{2}}=u_{1}$. The corresponding inverse $h$-function is $h_{\Pi}^{-1}(u_{1},u_{2})=u_{1}$. It follows that $h_{\theta}^{-1}(w_{t},w_{t-1})=w_{t}$, so the updating equation becomes $U_{t}=h_{\phi}^{-1}(w_{t},U_{t-1}),w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1)$, which is the updating equation of the classical copula-based time series model (see Eq. 10). ∎
This result is expected, since the proposed MAGMAR-Copula model is a generalization of the classical copula-based time series model. This result also justifies, describing the copula-based time series model as MAGMAR$(1,0)$-Copula model in further discussions. Next we define the MAG-Copula model.
Definition 3 (MAG(1)-Copula model).
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula time series. If the AR-copula is set to be the independence copula, $C_{\phi}=\Pi$, the resulting model is the MAG(1)-Copula model.
We note again that the MAG$(1)$-Copula model coincides with the $1$-dependent model from Chapter 3.14 of Joe (2014). It is easy to see that that a MAG(1)-Copula time series, $\{V_{t}\}_{t\in\mathbb{Z}}$ follows the updating equation
$V_{t}=h_{\theta}^{-1}(w_{t},w_{t-1}),\ w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1).$
Now we establish the stationarity of the MAG(1)-Copula model and show that the unconditional marginal distribution is given as standard uniform.
Proposition 3.
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAG$(1)$ time series. Then $\{U_{t}\}_{t\in\mathbb{Z}}$ is strictly stationary and $U_{t}\sim U(0,1),\ \forall t\in\mathbb{Z}$.
Proof.
See Appendix C
The models distributional properties are important for further work. They play an important role in establishing the stationarity of the MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula model with the normal AR-Copula, see Sect. 3.3.1. The next proposition treats the trivial case, where both, the AR and the MAG copula are chosen as the independence copula.
Proposition 4.
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula time series. If the AR and the MAG-copula are set to be the independence copula, $C_{\theta}=C_{\phi}=\Pi$ the time series $\{U_{t}\}_{t\in\mathbb{Z}}$ is an $iid$ standard uniform time series.
Proof.
Again we use that the inverse $h$-function associated to the independence copula is $h_{\Pi}^{-1}(u_{1},u_{2})=u_{1}$. From this it follows that the model updating equation is $U_{t}=w_{t}$, where $w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1)$. ∎
3.3 Distributional Properties of the model
In this section, the distributional properties of random variables $\{U_{i}\}_{i\in\mathbb{Z}}$, following the MAGMAR-Copula model, are derived. In order to keep the findings comprehensible, the ideas are developed for the MAGMAR(1,1)11(1,1)( 1 , 1 ) case. They can be generalized to the MAGMAR($p,q$) case.
First, we show that process is stationary for the case where the AR-copula is the normal copula and the parameter fulfills $|\phi|<1$. Second, we derive the conditional distribution of $U_{t}$ given past obervations. Third, using the law of total probability and conditional independence we derive the joint distribution of $\{U_{i}\}_{i\in\mathbb{Z}}$.
3.3.1 Stationarity
We show that the process is stationary for the case where the AR-copula is the normal copula. In order to do so, we consider the MAG representation of the MAGMAR$(1,0)$-normal-Copula time series and show that it does converge almost surely for all $t$. Together with the stationarity of the MAG$(1)$-Copula model, the stationarity of the MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula model will follow.
Proposition 5.
Let $\{U_{t}\}$ be a MAGMAR$(1,0)$-Copula time series where the AR-Copula is the normal copula, $C_{\phi}=C_{\phi}^{\text{normal}}$, with parameter $|\phi|<1$. Then the MAG()($\infty$)( ∞ ) representation,
$A_{t,n}:=\bigcirc_{i=0}^{n}h_{\phi}^{-1}(w_{t-i},\circ)$
does converge almost surely as $n\rightarrow\infty$ and $\lim_{n\rightarrow\infty}A_{t,n}=U_{t},\ U_{t}\sim U(0,1)$ almost surely.
Proof.
The inverse $h$-function associated to a normal copula with parameter $\phi$ is given as $h_{\phi}^{-1}(u_{1},u_{2})=\varphi(\varphi^{-1}(u_{1})(1-\phi^{2})^{\frac{1}{2}}+\phi\varphi^{-1}(u_{2}))$ (Smith et al., 2010). Then $A_{t,n}$ can be written as
$A_{t,n}=\varphi\left((1-\phi^{2})^{\frac{1}{2}}\sum_{i=0}^{n}\phi^{i}\varphi^{-1}(w_{t-i})\right).$
Since $w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U[0,1]$ and $\phi^{i}$ is decreasing geometrically in $i$, the variance of the sum $\sum_{i=0}^{n}\phi^{i}\varphi^{-1}(w_{t-i})$ is bounded and it converges almost surely (Billingsley, 1995). The limiting random variable is distributed as $N(0,\frac{1}{1-\phi^{2}})$. By the continuous mapping theorem and the continuity of $\varphi$ as well as the continuity of the multiplication with $(1-\phi^{2})^{\frac{1}{2}}$ it follows that $A_{t,n}$ also converges almost surely. Since $(1-\phi^{2})^{\frac{1}{2}}\sum_{i=0}^{n}\phi^{i}\varphi^{-1}(w_{t-i})\sim N(0,1)$, the limiting random variable is standard uniformly distributed. ∎
Notice that the same result holds when instead of an $iid$ sequence $w_{t}$, a dependent, but stationary, sequence is plugged in. This is in line with Theorem 36.4 from Billingsley (1995). Furthermore, from Lemma 5 it follows that the MAGMAR$(1,0)$-Copula time series is strictly stationary. We want to note, that the MAG($\infty$)-representation of the MAGMAR$(1,0)$-Copula time series is analogous to the MA($\infty$) representation in the classical linear ARMA model (see Hamilton, 2020) and is equivalent to the representation derived in McNeil and Bladt (2022). In principle the MA($\infty$) representation of the classical linear ARMA model can also be written as a infinite composition. However, there the composed functions are linear and, hence, also their composition is linear. We noted earlier that the $a.s.$ convergence in proposition 5 also holds if the innovation time series $\{w_{t}\}_{t\in\mathbb{Z}}$ is not $iid$ but merely stationary. In the next proposition we exploit this fact to show that the MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula time series with the AR-Copula being the normal copula is stationay.
Proposition 6.
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula time series where the AR-Copula is the normal copula, $C_{\phi}=C_{\phi}^{\text{normal}}$, with parameter $|\phi|<1$. Then $\{U_{t}\}_{t\in\mathbb{Z}}$ is strictly stationary.
Proof.
Similar to the reasoning in the proof of Lemma 5, the MAG()($\infty$)( ∞ ) representation of $\{U_{t}\}$ can be written as
$A_{t,n}=\varphi\left((1-\phi^{2})^{\frac{1}{2}}\sum_{i=0}^{n}\phi^{i}\varphi^{-1}(h_{\theta}^{-1}(w_{t-i},w_{t-i-1}))\right).$
Now notice that $\{V_{t},\ V_{t}:=h_{\theta}^{-1}(w_{t},w_{t-1})\}_{t\in\mathbb{Z}}$ is a MAG$(1)$ time series. From proposition 3 it follows that $\{V_{t}\}_{t\in\mathbb{Z}}$ is strictly stationary, so $\{\varphi^{-1}(V_{t})\}_{t\in\mathbb{Z}}$ is also strictly stationary. Since by proposition 5 converges almost surely and thereby constitutes a measurable mapping, by Theorem 36.4 of Billingsley (1995) it follows that the process is strictly stationary. ∎
Note that the stationary distribution of $\{U_{t}\}_{t\in\mathbb{Z}}$ does not necessarily have to be uniform. This may be problematic for forecasting and general modeling. A possible solution would be to apply the PIT to the random variables on the pseudo-observation scale before applying the quantile transformation. We would like to explore this in future research. Next, we derive the conditional distribution of $U_{t}$ given all past observations.
3.3.2 Conditional distribution
The conditional distribution of $U_{t}|\mathcal{F}_{t-1}$, where $\mathcal{F}_{s}:=\sigma(\{X_{r}\}_{r\leq s})$, can be derived by using the density transformation theorem (see e.g. Devore et al., 2012, Sect. 4.7). We can interpret the updating equation of the MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula model conditional on $(U_{t-i})_{i=1}^{\infty}:=(U_{t},U_{t-1},\ldots)$ as follows: The random variable $w_{t}\sim U(0,1)$ is transformed by a function $g:(0,1)\rightarrow(0,1)$ to the random variable $U_{t}$, so $U_{t}|\mathcal{F}_{t-1}=g(w_{t})|\mathcal{F}_{t-1}$. The other random variables, $(U_{t-i})_{i=1}^{\infty}$, are regarded as deterministic due to conditioning. In the function $g$, $(U_{t-i})_{i=1}^{\infty}$ are regarded as constants. We give an explicit form for $g$ in the proof of the following proposition.
Proposition 7.
Let $\{U_{t}\}_{t\in\mathbb{Z}}$ be a MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula time series. Furthermore, let the AR-Copula, $C_{\phi}$, and the MAG-Copula, $C_{\theta}$, be differentiable. Then the conditional density of $U_{t}$ given $\mathcal{F}_{t-1}$ is given as
$f_{U_{t}|\mathcal{F}_{t-1}}(x)=c_{\theta}\left(h_{\phi}(x,U_{t-1}),w_{t-1}(\{U_{t-i}\}_{i=1}^{\infty})\right)c_{\phi}(x,U_{t-1}).$ (17)
where $c_{\theta}$ is the copula density associated to the MAG-Copula and $c_{\phi}$ is the density associated to the AR-Copula. The expression $w_{t-1}((U_{t-i})_{i=1}^{\infty})$ is to be understood as a function of conditioning variables $(U_{t-i})_{i=1}^{\infty}$, which are regarded as constant in the conditional density.
Proof.
Let $g:(0,1)\rightarrow(0,1)$ be given as
$g(x)=$ $h_{\phi}^{-1}\left(h_{\theta}^{-1}(x,w_{t-1}((U_{t-i})_{i=1}^{\infty}),U_{t-1}\right).$
where $w_{t-1}((U_{t-i})_{i=1}^{\infty})$ is again a function of the conditioning variables $(U_{t-i})_{i=1}^{\infty}$. The explicit form of $w_{t}((U_{t-i})_{i=1}^{\infty})$ is given by the AR($\infty$) representation, Eq. 14. It holds that $U_{t}|\mathcal{F}_{t-1}=g(w_{t})|\mathcal{F}_{t-1}$. Application of the density transformation theorem (Devore et al., 2012, Sect. 4.7) yields the result. (The calculation of the Jacobian determinant is given in Appendix B) ∎
It is evident from the conditional density, Eq. 17, that $U_{t}|\mathcal{F}_{t-1}=U_{t}|(U_{t-i})_{i=1}^{\infty}=U_{t}|(U_{t-1},w_{t-1})$. This means that conditioning on all past random variables $(U_{t-i})_{i=1}^{\infty}$ is equivalent to conditioning on $(U_{t-1},w_{t-1})$. This is the same property as in the classical linear ARMA model. Note that for the special case where $c_{\theta}$ is the independence copula density, $c_{\theta}=1_{(0,1)^{2}}$, the conditional density of the copula-based time series model with autoregressive order of one is recovered. In the generalized MAGMAR($p,q$) case, the conditional density can be derived in a similar fashion. The random variable $w_{t}$ is transformed to the random variable $U_{t}$ by an appropriate transformation function $g$. The resulting density is a rather complicated expression which is given in Appendix E. Furthermore we note that multivariate conditional densities such as $f_{U_{t},U_{t-1}|\mathcal{F}_{t-2}}$ can be derived in a similar way. The derivation of joint conditional densities for the MAGMAR(1,1)11(1,1)( 1 , 1 ) model can be found in Appendix F.
3.3.3 Joint distribution/Likelihood
To derive the joint distribution of $t$ consecutive random variables from the time series $\{U_{t}\}_{t=1}^{T}$ following the MAGMAR-Copula model, conditional independence and the law of total probability (Schervish, 2012, B.3.5) is utilized. The joint distribution of the $t$ random variables $(U_{t},\ldots,U_{1})$ is given as
$f_{U_{t},\ldots,U_{1}}(u_{t},\ldots,u_{1})=f_{U_{1}}(u_{1})\prod_{i=0}^{t-2}f_{U_{t-i}|\mathcal{F}_{t-i-1}}(u_{t-i}|\mathcal{F}_{t-i-1}).$ (18)
[…]
3.4 Estimation & Modeling
In this section we describe how a univariate, strictly stationary time series may be modeled with the MAGMAR-Copula model. First, we describe the (pseudo-)maximum likelihood estimation given fixed AR and MAG orders and fixed bivariate copulas. Second, we describe challenges in model selection.
Given a fixed AR order $p$, fixed MA order $q$, fixed bivariate copula specifications $C_{\phi_{1}},\ldots,C_{\phi_{p}}$ and $C_{\theta_{1}},\ldots,C_{\theta_{q}}$ for the respective D-vines and observations $(u_{T},\ldots,u_{1})$, the estimation can be performed by (pseudo)-maximum likelihood estimation. The likelihood can be derived directly from the joint density (Eq. 18). The pseudo-likelihood of the MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula model of $T$ observations is given by
$L((\phi,\theta);u_{T},u_{T-1},\ldots,u_{1})=$ $\prod_{i=0}^{T-1}c_{\theta}[h_{\phi}(u_{T-i},u_{T-i-1}),w_{T-i-1}((u_{T-j-1})_{j=0}^{T-2})]$
×$\times c_{\phi}[u_{T-i},u_{T-i-1}].$ (19)
The likelihood of the MAGMAR$(p,q)$-Copula model is given in a completely analogous way using the Eq. 21 from Appendix E for the conditional densities. As in the ARMA and the GARCH model, the likelihood has to be calculated using an iterative scheme. The algorithm for calculating the likelihood of a MAGMAR$(p,q)$-Copula time series given AR & MAG orders, $p$ $q$, fixed bivariate copula specifications, $C_{\phi_{1}},\ldots,C_{\phi_{p}}$ and $C_{\theta_{1}},\ldots,C_{\theta_{q}}$ for the respective D-vines, fixed parameters $(\phi_{1},\ldots,\phi_{p},\theta_{1},\ldots,\theta_{q})$ and observations $(u_{T},\ldots,u_{1})$ is as follows.
  • 1)
    Set initial values for the first $s=\max\{p,q\}$ observations. (A natural choice for initial values in this context is 0.5 because for symmetric marginal distributions this corresponds to the expectation value. This is also the choice of initial values that we choose for our empirical application.)
  • 2)
    For $i\in\{s+1,\ldots,T\}$ calculate $R_{\phi_{1},\ldots,\phi_{p}}^{1}((u_{i-1},\ldots,u_{i-p}),u_{i})$ iteratively.
  • 3)
    For $i\in\{s+1,\ldots,T\}$ calculate
    $w_{i}=(R_{\theta_{1},\ldots,\theta_{q}}^{1})((w_{i-1},\ldots,w_{i-q}),R_{\phi_{1},\ldots,\phi_{p}}^{1}((u_{i-1},\ldots,u_{i-p}),u_{i}))$
    iteratively.
  • 4)
    For $i\in\{s+1,\ldots,T\}$ calculate $f_{U_{i}|\mathcal{F}_{i-1}}(u_{i})$ (Eq. 21 for MAGMAR$(p,q)$ and Eq. 17 for MAGMAR(1,1)11(1,1)( 1 , 1 )).
  • 5)
    Take the product of all conditional densities for the likelihood or the negative sum of log-densities for the negative log-likelihood: $L=\prod_{i=s+1}^{T}f_{U_{i}|\mathcal{F}_{i-1}}(u_{i})$ and $\text{nLL}=-\sum_{i=s+1}^{T}\log(f_{U_{i}|\mathcal{F}_{i-1}}(u_{i}))$, respectively.
Now we turn our attention to challenges in model selection. In contrast to linear models (i.e. ARMA), model selection in the MAGMAR-Copula context does not only amount to selecting appropriate model orders, i.e. AR & MAG orders, but also appropriate copulas. In popular vine copula software such as the VineCopula package in R (Schepsmeier et al., 2015), 38 bivariate copula specifications are available. Hence estimating all possible models is computationally expensive. In order to be able to perform sensible model selection, we suggest to be not solely relying on information criteria but also doing pre-selection based on descriptive statistics i.e. pair plots, acf, pacf or kpacf (see McNeil and Bladt, 2022, for the kpacf) and tail dependencies. The acf, pacf and kpacf can give information about sensible lag orders. When the partial dependence measures are cutoff after some lag order, a MAG part might not be necessary. When the partial auto-dependence measures are not cut-off, a long auto-dependence is indicated hence requiring inclusion of a MAG part in the model. The tail dependence measures can give information about sensible bivariate copula choices. No tail dependence indicates linear relations, while symmetric tail dependence indicates heavy-tailed dependence. Asymmetric tail dependence naturally indicates asymmetric copulas.
For further use, we introduce new notation. First, we abbreviate the t,normal,gumbel and independence copulas by ’t’, ’n’, ’g’ and ’i’, respectively. By the following example, we explain the new notation: We let MAGMAR($3,2$)-ngt-it denote the MAGMAR-Copula model with 3 AR-copulas and 2 MAG-copulas. The first AR-copula, governing the dependence of $U_{t}$ and $U_{t-1}$, is the normal copula, indicated by ’n’. The second AR-copula is the gumbel copula indicated by ’g’ and the third AR-copula is the t-copula, indicated by ’t’. The first MAG-copula is the independence copula, indicated by ’i’ and the second MAG-copula is the t-copula, indicated by ’t’. In the next section we model the US inflation with the MAGMAR-Copula model using the likelihood derived in this section.
4 Application to US Inflation
In this section, the modeling performance of the MAGMAR-Copula model is investigated. We compare its performance with the performance of the copula-based time series model with infinite partial dependence proposed by McNeil and Bladt (2022). Just as in McNeil and Bladt (2022), quarterly US inflation is be modeled. We use the consumer price index data, cpi from the tscopula package provided by the authors, and calculate the inflation as the growth rates of the CPI. The time series comprises 244 observations spanning from 1960Q1 to 2020Q4. Before estimating the respective models, the time series is transformed to pseudo-observations by applying the rescaled empirical cumulative distribution function. The models’ performances are assessed with the AIC and BIC. The results can be found in Table 1.
Table 1: Number of Parameters, negative log-likelihood and Information criterion values of the mentioned models fitted to quarterly US inflation data (244 observations). The MB-gaussian and the MB-gumbel model are the models from McNeil and Bladt (2022). ’gaussian’ and ’gumbel’ indicate which copula is chosen for the partial copulas. The results can also be found in their paper and can be recreated using the package provided by the authors. The MAGMAR$(p,q)$ model is the proposed autoregressive moving aggregate copula-based time series model with $p$ autoregressive and $q$ moving aggregate orders. The letters after the model orders indicate which copulas are used at the respective order, as is described in Sect. 3.4. The highlighted values are the lowest ones in the respective column.
Model # pars nLL AIC BIC
MB-normal $6$ $-98.31$ $-184.62$ $-163.64$
MB-gumbel $6$ $-110.64$ $-209.28$ $-188.30$
MAGMAR(4, 0)-ggtg $5$ $-111.05$ $-212.10$ $-194.61$
MAGMAR(4, 1)-nnnn-n $5$ $-93.48$ $-176.95$ $-159.47$
MAGMAR(4, 1)-gggg-t $6$ $-110.88$ $-209.77$ $-188.79$
MAGMAR(4, 1)-ggtg-t $7$ $-113.60$ $-213.21$ $-188.72$
MAGMAR(4, 1)-ging-t $5$ $-112.16$ $\mathbf{-214.31}$ $\mathbf{-196.83}$
The MAGMAR(4,1)-ging-t model minimizes the AIC, as well as the BIC. The necessity of including the gumbel copula indicates asymmetric temporal dependence while the inclusion of the t-copula indicates heavy-tailedness of the dependence structure. All of the best performing models incorporate 4 autoregressive orders. This result indicates seasonality in the quarterly inflation. The necessity of including a MAG part indicates that the data generating process of US inflation may not fulfill the Markov assumption. Our proposed model minimizes the AIC/BIC in comparison with the copula-based time series model with infinite partial dependence by McNeil and Bladt (2022). However, the model by McNeil and Bladt (2022) is still competitive. Now the reader may wonder which model to use in further applications. In this section we want to give our thoughts regarding this question. One advantage of the copula-ARMA model in comparison to our model is that the partial dependence structures, $C_{U_{t},U_{t-k}|U_{t-1},\ldots,U_{t-k+1}},\ k\in\mathbb{N}$, are explicitly accessible. In the MAGMAR-Copula model, the partial dependence structures are generally not explicitly accessible. To calculate e.g. $C_{U_{t},U_{t-1}}$ in the MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula model, first, the joint distribution of $(w_{t},U_{t})$ has to be calculated explicitly. This amounts to solving a homogeneous Fredholm integral equation of second kind (compare Atkinson, 1967, Micula and Milovanović, 2023). Having found the joint distribution of $(w_{t},U_{t})$, it can be used to calculate the partial dependencies by integration with respect to the respective measure or by density transformation. In general the equations may not always be solvable. However, we want to note that for the MAGMAR-Copula model conditional dependence structures, e.g. $f_{U_{t},U_{t-1}|\mathcal{F}_{t-2}}$ can be calculated explicitly, see Appendix F.
One major disadvantage of the copula-ARMA model is that it is not easy to incorporate two-parametric copulas, while our proposed model is able to incorporate them easily. The inclusion of two parametric copulas, especially the t-copula can be important when dealing with volatile time series. The volatility of the time series corresponds to a heavy-tailed temporal dependence structure (Loaiza-Maya et al., 2018, Pappert and Arsova, 2023). Hence, we think that our proposed model would be more suitable to model such time series.
5 Conclusion
Non-linear temporal dependencies as well as deviations from normal or elliptical distributions are apparent in many univariate time series. One approach to capture these effects is to employ a copula $C_{\phi}$ for the temporal dependence structure equipped with a suitable marginal distribution $F_{X}$. The resulting model is the well-known copula-based time series model. Alongside stationarity, the model assumes a finite Markov order $p$. This means that the distributional properties of a random variable of the time series, $X_{t}$, can only be conditionally influenced by the preceding $p$ values, $X_{t-1},\ldots,X_{t-p}$. This results in the model not being able to properly capture long autoregressive effects, i.e. persistency. Properly modeling persistency in the mean and the variance was a major improvement for ’classical’ time series modeling. To this end, the moving average term (MA) was introduced to autoregressive (AR) models resulting in the ARMA model. The generalization of the autoregressive conditional heteroskedasticity models (ARCH) was introduced to account for persistency in the variance, i.e. volatility clustering. Inspired by this approach we introduce a moving aggregate (MAG) part to the model equation of copula-based time series models. The moving aggregate is a non-linear function of the current and past innovations. Its functional form is given by a conditional copula. The resulting model is the Moving Aggregate Modified Autoregressive Copula-Based Time Series Model (MAGMAR-Copula time series model). After introducing the model we discuss the model components and nested models. The MAGMAR-Copula model can be understood as a generalization to classical copula-based time series models and the classical ARMA model since both models are nested. In comparison with the copula-ARMA model from McNeil and Bladt (2022) our model is in general not able to give explicit forms for partial dependencies, while the copula-ARMA model is able to do so. The main advantage of our model is that it is able to incorporate copulas with an arbitrary amount of parameters without any effort. This is especially relevant for the $t$-copula and related copulas. We derive both the MAG($\infty$) and the AR($\infty$) representations and continue to derive distributional properties such as the conditional marginal distribution and the likelihood function. Furthermore, we prove that the MAGMAR-Copula model is stationary for the case where the AR-copula is the normal copula and the parameter fulfills $|\phi|<1$. In an application to US inflation we compare our model with the model for infinite order partial dependence of McNeil and Bladt (2022). In their paper the authors introduce the copula-ARMA model, which is also able to capture the infinite memory of a time series. The comparison in the application is conducted by using information criteria (AIC and BIC). We find that our model is competitive with the model from McNeil and Bladt (2022).
Although our proposed MAGMAR-Copula model is quite general, it is only a model for univariate time series. The incorporation of other variables as covariates is, in principle, possible, but may not necessarily enhance the performance of the model in its current form. Modifications to properly account for multivariate effects is a question for future research. One possibility to extend our model to be multivariate is to consider the popular copula-GARCH model and employ MAGMAR-Copulas for temporal modeling instead of GARCH (Hu (2006), Jondeau and Rockinger (2006), Aloui et al. (2014) are standard references for the copula-GARCH model while Berrisch et al. (2022) is a recent application). Furthermore, we would like to explore the forecasting performance of our model. As mentioned earlier, our model is able to capture volatility clusters. It will be interesting to see how our model performs in comparison with ARCH-models. From a theoretical point of view, further distributional properties are of interest. Especially interesting are the joint distributions of consecutive random variables in the time series, $U_{t}$ and $U_{t-1}$. This joint distribution is related to the stationary distribution of the time series, which is also of high interest.
Acknowledgement
This research was (partially) funded in the course of TRR 391 Spatio-temporal Statistics for the Transition of Energy and Transport (520388526) by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation).
Appendix A Recovering classical linear ARMA
The classical linear ARMA(1,1) model can be recovered from equation Eq. 13. Given random variables $\{X_{t}\}_{t\in\mathbb{Z}}$, following the MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula model with marginal distribution function $F$,
$h_{\alpha}(F(X_{t}),F(X_{t-1}))=h^{-1}_{\beta}(w_{t},w_{t-1}),\ \ w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1),$
the linear ARMA model with AR parameter $\alpha$, MA parameter $\frac{\beta}{\sqrt{1-\beta^{2}}}$ and variance $(1-\alpha^{2})(1-\beta^{2})$ can be recovered when setting the copulas to the normal copulas and the marginal distribution $F$ to the standard normal distribution. It is well known in the literature that the AR model can be recovered from the copula-based time series model without MAG modification, see e.g. Chen and Fan (2006) and Smith et al. (2010). For completeness we also give the derivation here.
Proof.
The explicit forms for the $h$-functions are taken from Smith et al. (2010):
$h_{\alpha}^{\text{normal}}(u_{1},u_{2})=$ $\varphi([\varphi^{-1}(u_{1})-\alpha\varphi^{-1}(u_{2})](1-\alpha^{2})^{-\frac{1}{2}})$
$(h_{\alpha}^{\text{normal}})^{-1}(u_{1},u_{2})=$ $\varphi(\varphi^{-1}(u_{1})(1-\alpha^{2})^{\frac{1}{2}}+\alpha\varphi^{-1}(u_{2}))$
Where $\phi$ is the cdf of the standard normal distribution. Setting $C_{\alpha}=C_{\alpha}^{\text{normal}}$, $C_{\beta}=C_{\beta}^{\text{normal}}$ with $|\alpha|<1,|\beta|<1$ and $F_{X}=\varphi$ we get
[…]
$=$ $\varphi([\varphi^{-1}(\varphi(X_{t}))-\alpha\varphi^{-1}(\varphi(X_{t-1}))](1-\alpha^{2})^{-\frac{1}{2}})$
$=$ $\varphi([X_{t}-\alpha X_{t-1}](1-\alpha^{2})^{-\frac{1}{2}})$
[…]
Setting LHS=RHS and setting $\varphi^{-1}(w_{t})=:\varepsilon_{t},\ \forall t\in\{1,\ldots,T\}$, where due to the quantile transformation $\varepsilon_{t}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$, we obtain
(((( $X_{t}-\alpha X_{t-1})(1-\alpha^{2})^{-\frac{1}{2}}=\varepsilon_{t}(1-\beta^{2})^{\frac{1}{2}}+\beta\varepsilon_{t-1},$
$\iff$ $X_{t}-\alpha X_{t-1}=(1-\alpha^{2})^{\frac{1}{2}}(1-\beta^{2})^{\frac{1}{2}}\varepsilon_{t}+\beta(1-\alpha^{2})^{\frac{1}{2}}\varepsilon_{t-1},$
$\iff$ $X_{t}-\alpha X_{t-1}=\tilde{\varepsilon}_{t}+\frac{\beta}{(1-\beta^{2})^{\frac{1}{2}}}\tilde{\varepsilon}_{t-1}.$
This is the model equation for the ARMA($1,1$) model with AR-parameter $\alpha$, MA-parameter $\frac{\beta}{(1-\beta^{2})^{\frac{1}{2}}}$ and white noise normal innovations with variance $(1-\alpha^{2})(1-\beta)^{2}$. ∎
Appendix B Conditional distribution MAGMAR(1,1)
The conditional distribution of $U_{t}|\mathcal{F}_{t-1}$ in the MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula model can be derived by using the density transformation (see Devore et al., 2012, Sect. 4.7). The random variable $w_{t}$ is transformed by the function $g$ to the random variable $U_{t}$. The appropriate choice of $g$ is given as,
$g(x)=$ $h_{\phi}^{-1}(h_{\theta}^{-1}(x,w_{t-1}),U_{t-1})\ \ \ \text{where}\ U_{t-1},w_{t-1}\ \text{deterministic},$
$\Rightarrow g^{-1}(x)=$ $h_{\theta}(h_{\phi}(x,U_{t-1}),w_{t-1})),$
$\Rightarrow|\frac{d}{dx}g^{-1}(x)|=$ $c_{\theta}(h_{\phi}(x,U_{t-1}),w_{t-1})c_{\phi}(x,U_{t-1}).$
Using that $w_{t}\sim U(0,1)$ and hence $f_{w_{t}}(x)=1_{(0,1)}(x)$, the resulting conditional density for $U_{t}|\mathcal{F}_{t-1}$ is
$f_{U_{t}|\mathcal{F}_{t-1}}(x)=c_{\theta}(h_{\phi}(x,U_{t-1}),w_{t-1})c_{\phi}(x,U_{t-1}).$
Appendix C Distributional Properties of MAG(1)
The MAG(1) model is given by
$U_{t}=h_{\theta}^{-1}(w_{t},w_{t-1}),\ \ w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1)$
The marginal distribution is standard uniform and the time series is stationary.
Proof.
For $a\in(0,1)$, it holds that
$\mathbb{P}(U_{t}\leq a)=$ $\mathbb{P}(h_{\theta}^{-1}(w_{t},w_{t-1})\leq a),$
$=$ $\mathbb{E}(h_{\theta}(a,w_{t-1})),$
[…]
For $a\leq 0,\ \mathbb{P}(U_{t}\leq a)=0$ and for $a\geq 1,\ \mathbb{P}(U_{t}\leq a)=1$ since $U_{t}\in(0,1)$. So $\mathbb{P}(U_{t}\leq a)$ is the distribution function of the uniform distribution. The joint distribution of $(U_{t},U_{t-1})$ can be calculated as follows
$F_{U_{t},U_{t-1}}(a,b)=$ $\mathbb{P}(U_{t}\leq a,U_{t-1}\leq b)$
$=$ $\mathbb{P}(h_{\theta}^{-1}(w_{t},w_{t-1})\leq a,h_{\theta}^{-1}(w_{t-1},w_{t-2})\leq b)$
$=$ $\mathbb{E}(h_{\theta}(a,w_{t-1})1_{\{w_{t-1}\leq h_{\theta}(b,w_{t-2}\}})$
[…]
Note that implicitly it was used that the marginal distribution is of $U_{t}$ is standard uniform and time independent. Now using that the joint distribution of $(U_{t},U_{t-1})$ is time independent, it can be shown that also $F_{U_{t},U_{t-1},U_{t-2}}$ is time independent. This procedure can be generalized and then by induction it follows that the process is strictly stationary. (Furthermore it can be seen that $F_{U_{t},U_{t-1}}(a,b)$ is increasing in $a$ and $b$, the triangle inequality is fulfilled and $F_{U_{t},U_{t-1}}(1,b)=b$ as well as $F_{U_{t},U_{t-1}}(a,1)=a$. Hence $F_{U_{t},U_{t-1}}(a,b)$ is a copula.) ∎
Appendix D AR and MAG representation of the MAGMAR(p,q) model
The AR representation of the MAGMAR$(p,q)$-model (Eq. 16) can be derived by substituting for all $\{w_{t-i}\}_{i=1}^{\infty}$ iteratively. We start with the following expression:
$w_{t}=R_{\theta_{1},\ldots,\theta_{q}}^{1}\left((w_{t-1},\ldots,w_{t-q}),R_{\phi_{1},\ldots,\phi_{p}}((U_{t-1},\ldots,U_{t-p}),U_{t})\right).$
Now, $w_{t-1},\ldots,w_{t-q}$ have to be substituted again. Formally this amount to composing the function $R_{\theta_{1},\ldots,\theta_{q}}^{1}(\circ,R_{\phi_{1},\ldots,\phi_{p}}^{1}((U_{t-1},\ldots,U_{t-p}),U_{t}):(0,1)^{q}\rightarrow(0,1)$ with a vector of functions:
$w_{t}=$ $R_{\theta_{1},\ldots,\theta_{q}}^{1}(\circ,R_{\phi_{1},\ldots,\phi_{p}}^{1}((U_{t-1},\ldots,U_{t-p}),U_{t})$
$\circ\begin{pmatrix}R_{\theta_{1},\ldots,\theta_{q}}^{1}(\circ,R_{\phi_{1},\ldots,\phi_{p}}^{1}((U_{t-2},\ldots,U_{t-p-1}),U_{t-1})\\ \vdots\\ R_{\theta_{1},\ldots,\theta_{q}}^{1}(\circ,R_{\phi_{1},\ldots,\phi_{p}}^{1}((U_{t-q-1},\ldots,U_{t-p-q}),U_{t-q})\end{pmatrix}\circ\ldots$
Now the $\{w_{t-i}\}_{i=2}^{\infty}$ in each vector has to be substituted for again. The input for each element of the vector is again a $q$-dimensional vector. The total dimension of the input vector at the third level hence is given by $q^{2}$. We shorten notation by introducing the $(\stackrel{{\scriptstyle\rightarrow}}{{\bigcirc}}_{d})_{i\in\mathbb{N}_{0}}$ symbol which acts as follows for a sequence of functions $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, $i\in\mathbb{N}_{0}$
$(\stackrel{{\scriptstyle\rightarrow}}{{\bigcirc}}_{d})_{i\in\mathbb{N}_{0}}f_{i}(\circ)=f_{0}\circ\begin{pmatrix}f_{1}\\ \vdots\\ f_{d}\end{pmatrix}\circ\begin{pmatrix}f_{d+1}\\ \vdots\\ \vdots\\ f_{d+d^{2}}\end{pmatrix}\circ\ldots$
Using the newly introduced notation, the AR representation may be written as
$w_{t}=(\stackrel{{\scriptstyle\rightarrow}}{{\bigcirc}}_{q})_{i\in\mathbb{N}_{0}}R^{1}_{\theta_{1},\ldots,\theta_{q}}(\circ,R_{\phi_{1},\ldots,\phi_{p}}((U_{t-1-i},\ldots,U_{t-p-i}),U_{t-i})$ (20)
Using the same notation, the MAG($\infty$)-representation can be written as
$U_{t}=$ $R_{\phi_{1},\ldots,\phi_{p}}^{-1}(\circ,R_{\theta_{1},\ldots,\theta_{q}}^{-1}((w_{t-1},\ldots,w_{t-q}),w_{t}))$
$\circ\begin{pmatrix}R_{\phi_{1},\ldots,\phi_{p}}^{-1}(\circ,R_{\theta_{1},\ldots,\theta_{q}}^{-1}((w_{t-2},\ldots,w_{t-q-1}),w_{t-1}))\\ \vdots\\ R_{\phi_{1},\ldots,\phi_{p}}^{-1}(\circ,R_{\theta_{1},\ldots,\theta_{q}}^{-1}((w_{t-p-1},\ldots,w_{t-p-q}),w_{t-p}))\end{pmatrix}\circ\ldots$
$=$ $(\stackrel{{\scriptstyle\rightarrow}}{{\bigcirc}}_{p})_{i\in\mathbb{N}_{0}}R_{\phi_{1},\ldots,\phi_{p}}^{-1}(\circ,R_{\theta_{1},\ldots,\theta_{q}}^{-1}((w_{t-1-i},\ldots,w_{t-q-i}),w_{t-i}))$
Appendix E Conditional Distribution MAGMAR(p,q)
The conditional distribution of $U_{t}$ given all past observations is given by:
$f_{U_{t}|\mathcal{F}_{t-1}}(x)=$ $c_{\theta_{q}}\left[R_{\phi_{1},\ldots,\phi_{p}}^{1}((u_{t-1},\ldots,u_{t-p}),x),R^{1}_{\theta_{1},\ldots,\theta_{q-1}}((w_{t-2},\ldots,w_{t-q}),w_{t-1})\right]$
$\prod_{i=0}^{p-1}c_{\phi_{p-i}}\bigg{[}R^{1}_{\phi_{1},\ldots,\phi_{p-1-i}}((u_{t-1},\ldots,u_{t-p+1+i}),x),$
$R^{2}_{\phi_{1},\ldots,\phi_{p-1-i}}((u_{t-1},\ldots,u_{t-p+1+i}),u_{t-p+i})\bigg{]}.$ (21)
The proof is analog to the proof for the MAGMAR(1,1)-Copula case.
Appendix F Joint Conditional Distribution of the MAGMAR model
The joint conditional distribution of $(U_{t},U_{t-1})|\mathcal{F}_{t-2}$ is derived. The density transform theorem for multivariate transformations (see Devore et al., 2012, 5.6) is utilized. The two random variables $w_{t}$ and $w_{t-1}$ are transformed by the function $G$ to the random variables $U_{t}$ and $U_{t-1}$. The appropriate choice for $G$ and the corresponding inverse and Jacobian are given by,
$\begin{pmatrix}z\\ w\end{pmatrix}:=$ $G((a,b)^{T})=\begin{pmatrix}h_{\phi}^{-1}(h_{\theta}^{-1}(a,w_{t-2}),u_{t-2})\\ h_{\phi}^{-1}(h_{\theta}^{-1}(b,a),h_{\phi}^{-1}(h_{\theta}^{-1}(a,w_{t-2}),u_{t-2}))\end{pmatrix},$
$=$ $\begin{pmatrix}h_{\phi}^{-1}(h_{\theta}^{-1}(a,w_{t-2}),u_{t-2})\\ h_{\phi}^{-1}(h_{\theta}^{-1}(b,a),z)\end{pmatrix}.$
$\begin{pmatrix}a\\ b\end{pmatrix}:=$ $G^{-1}(z,w)=\begin{pmatrix}h_{\theta}(h_{\phi}(z,u_{t-2}),w_{t-2})\\ h_{\theta}(h_{\phi}(w,z),h_{\theta}(h_{\phi}(z,u_{t-2}),w_{t-2}))\end{pmatrix},$
$=$ $\begin{pmatrix}h_{\theta}(h_{\phi}(z,u_{t-2}),w_{t-2})\\ h_{\theta}(h_{\phi}(w,z),a)\end{pmatrix}$
$\mathcal{J}=$ $\begin{pmatrix}c_{\theta}[h_{\phi}(z,u_{t-2}),w_{t-2}]c_{\phi}[z,u_{t-2}]&0\\ \partial_{z}(\ldots)&c_{\theta}[h_{\phi}(w,z),h_{\theta}(h_{\phi}(z,u_{t-2}),w_{t-2})]c_{\phi}[w,z]\end{pmatrix},$
$|\mathcal{J}|=$ $c_{\theta}[h_{\phi}(z,u_{t-2}),w_{t-2}]c_{\phi}[z,u_{t-2}]c_{\theta}[h_{\phi}(w,z),h_{\theta}(h_{\phi}(z,u_{t-2}),w_{t-2})]c_{\phi}[w,z].$
Here $u_{t-2}$ and $w_{t-2}$ ought to be understood as (deterministic by conditioning) constants. Since $w_{t}$ and $w_{t-1}$ are independent and uniformly distributed their contribution to the conditional density is just $1_{(0,1)^{2}}$. The density can be formulated as
$f_{U_{t},U_{t-1}|\mathcal{F}_{t-2}}(z,w)=f_{w_{t},w_{t-1}}(z,w)|\mathcal{J}|$
$=$ $c_{\theta}[h_{\phi}(z,u_{t-2}),w_{t-2}]c_{\phi}[z,u_{t-2}]c_{\theta}[h_{\phi}(w,z),h_{\theta}(h_{\phi}(z,u_{t-2}),w_{t-2})]c_{\phi}[w,z].$
The partial conditional density of $U_{t},U_{t-k}|U_{t-1},\ldots,U_{t-k+1},U_{t-k-1},\ldots$ can again be derived by using the density transformation theorem. The random variables $w_{t}$ and $w_{t-k}$ are transformed to the random variales $U_{t}$ and $U_{t-k}$. The appropriate choice of the transformation $G$ is,
$G((a,b)^{T})=$ $\begin{pmatrix}z\\ w\end{pmatrix}=\begin{pmatrix}h_{\phi}^{-1}(h_{\theta}^{-1}(a,w_{t-k-1}),u_{t-k-1})\\ h_{\phi}^{-1}(h_{\theta}^{-1}(b,w_{t-1}(a)),u_{t-1})\end{pmatrix}$
$G^{-1}(z,w)=$ $\begin{pmatrix}a\\ b\end{pmatrix}=\begin{pmatrix}h_{\theta}(h_{\phi}(z,u_{t-k-1}),w_{t-k-1})\\ h_{\theta}(h_{\phi}(w,u_{t-1}),w_{t-1}(z))\end{pmatrix}$
$\mathcal{J}=$ $\begin{pmatrix}c_{\theta}[h_{\phi}(z,u_{t-k-1}),w_{t-k-1})c_{\phi}[z,u_{t-k-1}]&0\\ \partial_{z}(\ldots)&c_{\theta}[h_{\phi}(w,u_{t-1}),w_{t-1}(z)]c_{\phi}[w,u_{t-1}]\end{pmatrix}$
$|\mathcal{J}|=$ $c_{\theta}[h_{\phi}(z,u_{t-k-1}),w_{t-k-1})c_{\phi}[z,u_{t-k-1}]c_{\theta}[h_{\phi}(w,u_{t-1}),w_{t-1}(z)]c_{\phi}[w,u_{t-1}].$
Here $w_{t-1}(z)$ means evaluating the AR($k-1$) representation of $w_{t-1}$ at $u_{t-1},\ldots,u_{t-k+1}$ and $w_{t-k}=h_{\theta}(h_{\phi}(z,u_{t-k-1}),w_{t-k-1})$ and $U_{t-k}=z$.
Appendix G Copulas used in this Work
Normal Copula: The Normal copula models linear dependencies. It is constructed by the inversion method using the multivariate normal distribution (Smith, 2021, Nelsen, 2006),
$C^{\text{normal}}_{\Sigma}[u_{1},\ldots,u_{p}]=\Phi_{\Sigma}(\varphi^{-1}(u_{1}),\ldots,\varphi^{-1}(u_{d})].$ (22)
Here $\Phi_{\Sigma}$ denotes the multivariate normal distribution function with correlation matrix $\Sigma\in[-1,1]^{p\times p}$, symmetric and positive definite. $\varphi^{-1}$ denotes the quantile function of the univariate standard normal distribution.
$t$-copula: The $t$-copula models linear dependencies with heavy tails. Similar to the normal copula it is also constructed by the inversion method. It is given as (Demarta and McNeil, 2005),
$C^{t}_{\Sigma,\nu}[u_{1},\ldots,u_{p}]=t_{\Sigma,\nu}(t_{\nu}^{-1}(u_{1}),\ldots,t_{\nu}^{-1}(u_{p})),$ (23)
where $t_{\Sigma,\nu}$ denotes the multivariate t-distribution with correlation matrix $\Sigma\in[-1,1]^{p\times p}$, symmetric and positive definite, degrees of freedom $\nu>0$ and $t_{\nu}^{-1}$ is the quantile function of the standard $t$-distribution with degrees-of-freedom $\nu$. For $\nu\rightarrow\infty$, the $t$-copula approaches the normal copula. For small $\nu$, the dependence structure given by the $t$-copula is heavy-tailed, meaning that extreme events coincide.
Gumbel Copula: The gumbel copula is an archimedean copula (see Hofert, 2008, Genest and Rivest, 1993), hence it is constructed via $C[u_{1},\ldots,u_{p}]=\psi(\psi^{-1}(u_{1})+\ldots+\psi^{-1}(u_{p}))$ by choosing a continuous and non-increasing generator function $\psi$. The generator function for the gumbel copula is given as $\psi(t)=\exp(-t^{\frac{1}{\rho}})$ (Hofert, 2008). The resulting copula is
$C^{\text{gumbel}}_{\rho}=\exp(-((-\ln(u_{1}))^{\rho}+\ldots+(-\ln(u_{p}))^{\rho})^{\frac{1}{\rho}}).$ (24)
The gumbel copula has non-zero upper tail dependence while having no lower tail dependence (Nelsen, 2006, Ex. 5. 22.), hence displaying an asymmetric dependence structure.
Independence Copula: The independence copula simply models independence. Accordingly it is given by the product of its arguments (Nelsen, 2006),
$\Pi[u_{1},\ldots,u_{p}]=\prod_{i=1}^{p}u_{i}.$ (25)
Choosing $\Pi$ as the copula in Eq. 1, it can be seen, that the joint distribution is just the product of marginals. Hence, random variables having the independence copula as their copula implies independence and vice versa. For $\Sigma=I_{p\times p}$, the normal copula becomes the independence copula. For $\Sigma=I_{p\times p}$ and $\nu\rightarrow\infty$ the $t$-copula becomes the independence copula and for $\rho=1$, the gumbel copula becomes the independence copula.